A perfectly competitive firm has the cost function TC = 1000 + 2Q + 0.1 Q2. What is the lowest price at which this firm can break even?


1
Expert's answer
2021-05-24T10:53:39-0400

"TC = 1000+2Q+0.1Q^2 \\

MC = 2 + 0.2Q \\

AC = \frac{1000}{Q} + 2 + 0.1Q"

Set MC=AC for minimum AC

"2 + 0.2Q = \frac{1000}{Q} + 2 + 0.1Q \\

0.1Q = \frac{1000}{Q} \\

0.1Q^2 = 1000 \\

Q^2 = 10000 \\

Q = 100 \\

AC = \frac{1000}{100}+2+0.1 \times 100 \\"

= 10 + 2 +10 = $22

This is the lowest price at which the firm can break even.